A cylinder hold 400centmitres cube of juice when it is full.Determine the height of the cylinder in terms of the radius.Please send a massege with me.
Answer by Cromlix(4381)

Volume of cylinder = pi r^2 h
400 cm^3 = pi r^2 h
r^2h = 400/pi
h = 127.3/r^2
Hope this helps
:-)